OpenStudy (anonymous):

2y+3x=8 y=2x-3 graph the system

OpenStudy (anonymous):

We need to find x-intercept (where y=0) and y-intercept (where x=0) before we can graph. Can you find those intercepts?

hero (hero):

4x-6 + 3x = 8 7x-6 = 8 7x = 14 x = 2 y = 1

OpenStudy (anonymous):

Graphing the system is the problem statement.

hero (hero):

I know

OpenStudy (anonymous):

I know you know! :-)

OpenStudy (anonymous):

Please help out. I can't bring myself to give the answer.

hero (hero):

Sorry

OpenStudy (anonymous):

well when i put found the x and y intercepts of the first equation is x inter =2.66 when y=0 (2.66, 0) rite

hero (hero):

I'd rather just post the solution and move on.

OpenStudy (anonymous):

Correct. Or, (8/3,0). Y-intercept? Same for second equation?

OpenStudy (anonymous):